If a cord is under a tension of F T = 135 N, and has a mass per unit length of 0.10 kg/m, what power is

required to transmit a 120-Hz transverse wave of amplitude 2.0 cm?


1
Expert's answer
2021-10-24T18:23:18-0400

The power in transverse wave is given as follows (see https://opentextbc.ca/universityphysicsv1openstax/chapter/16-4-energy-and-power-of-a-wave/):


"P = \dfrac12\mu A^2\omega^2v"

where "\mu = 0.10kg/m", "A = 2.0cm = 0.020m",


"\omega = 2\pi f"

where "f = 120Hz" and


"v = \sqrt{\dfrac{F_T}{\mu}}"

where "F_T = 135N". Combining it all together, obtain:


"P = \dfrac12\mu A^2(2\pi f)^2\sqrt{\dfrac{F_T}{\mu}} = 2\pi^2\mu A^2f^2\sqrt{\dfrac{F_T}{\mu}}\\
P = 2\pi^2\cdot 0.1\cdot 0.02^2\cdot 120^2\cdot \sqrt{\dfrac{135}{0.1}} \approx 4.2\times 10^2W"

Answer. "4.2\times 10^2W"